abbotttsville and broken branch are distance 175 millas apart. on a map, the distance between the two towns are 2.5 inches. what is the scale of the map?
Found 2 solutions by rfer, Alan3354:
Answer by rfer(16322)

175/2.5=70 miles per inch

Answer by Alan3354(69443)

abbotttsville and broken branch are distance 175 millas apart. on a map, the distance between the two towns are 2.5 inches. what is the scale of the map?
--------------------
175/2.5 = 70 miles per inch
= 70*5280*12 inch/inch
= 1:4435200
--------------------
The tactical maps (tac maps) I used in Vietnam were 1:50000.
The maps used on the ground are usually 1:10000 or 1:25000.
High speed high altitude aircraft use 1:1,000,000 or 1:2,500,000.